1. 3y=2y-18
2. 5y=4y+10
3. 2y+15=y
4. 4y-26=3y
5. 7x-14=6x-8

Answer by ReadingBoosters(3246)

1)
3y = 2y-18
3y - 2y = -18
y = -18
...
2)
5y = 4y + 10
5y - 4y = 10
y = 10
...
3)
2y + 15 = y
2y - y = -15
y = -15
...
4)
4y - 26 = 3y
4y - 3y = 26
y = 26
...
5)
7x - 14 = 6x - 8
7x - 6x = -8 + 14
x = 6
.....................
Delighted to help.
-Reading Boosters
Website: www.MyHomeworkAnswers.com
Wanting for others what we want for ourselves.